A chemist mixes 10L of a solution that is 30% acid with 30L of a solution that is 50% acid. What is the percent of acid in the new mixture?
Answer by Maths68(1474)


Solution A
=========
Amount = 10ml
Concentration =30% =0.3
Solution B
=========
Amount = 30ml
Concentration =50% = 0.5
Mixture
========
Amount =10+30 = 40ml
Concentration =x
[Amount Solution A * Concentration A] + [Amount Solution B * Concentration of B] =[Amount of Mixture *Concentration of Mixture]
(10)(0.3)+(30)(0.5)=(40)(x)
3+15=40x
18=40x
x=18/40
x=0.45
x=45/100=45%
A chemist mixes 10L of a solution that is 30% acid with 30L of a solution that is 50% acid to prepare 40ml mixture that is 45% acid.